A high school student in Seattle, Washington, attends the University of Central Florida. On the way to UCF he took a southern route. After graduation he returned to Seattle via a northern trip. On both trips he had the same average speed. If the southern trek took 45 hours, the northern trek took 50 hours, the northern trek was 310 miles longer, how long was each trip?
The southern trip is miles long The northern trip is miles long

Answer by graphmatics(170)

Let S be the southern trip distance, then S+310 is the northern trip distance. Let x be the speed of the trips. S/x=45. (S+310)/x=50. S=45*X. S+310=50*X. 45*x+310=50*x. 5*x=310. x=62 So Southern trip distanceS=45*62=2790 miles. Northern trip distance=2790+310=3100 miles.